OpenStudy (gamer56):

Select the mathematical sentence that is true. A. 2 • (8 – 5) + 3 < 7 + 3 • 2 B. 2 • (8 – 5) + 3 > 7 + 3 • 2 C. 2 • (8 – 5) + 3 = 7 + 3 • 2

OpenStudy (gamer56):

@shamil98 @shamil98 @kitteh♥ @ztp5031

OpenStudy (anonymous):

C'mon mate, this is basic math.

OpenStudy (anonymous):

PEMDAS

OpenStudy (gamer56):

Im redoing some stuff can u tell me how or guide me in some way???

OpenStudy (anonymous):

8-5+3?

OpenStudy (gamer56):

8-5=3 3+3=6

OpenStudy (anonymous):

7+3×2?

OpenStudy (gamer56):

7+3=10 10 x 2=20

OpenStudy (anonymous):

is 6<20?

OpenStudy (gamer56):

yes 20 is bigger then 6 if thats what your asking

OpenStudy (anonymous):

ok so statement 1 is true

OpenStudy (gamer56):

ITS A !!!

OpenStudy (gamer56):

o

OpenStudy (gamer56):

can you help with 3 more?

OpenStudy (anonymous):

statement 2 is false because it's saying 6 is greater than 20 and statement 3 is false because it's saying 6 is equal to 20

OpenStudy (anonymous):

i can try

OpenStudy (gamer56):

Find the solution set for the given inequality if the replacement set for x is the set {1, 2, 3, 4}. 4x – 3 ≥ 5

OpenStudy (anonymous):

4(1)-3?

OpenStudy (gamer56):

Assuming that i do 4 x 1 = 4 - 3=1 Correct??

OpenStudy (anonymous):

yes always follow this rule Pemdas Parenthesis Exponents Multiply Divide Add Subtract

OpenStudy (gamer56):

:D

OpenStudy (anonymous):

is 1 greater than or equal to 5?

OpenStudy (gamer56):

neither 5 is greater then 1 correct?

OpenStudy (anonymous):

yes correct so 1 wont be in the function I'm in and out because i have to catch up in my math class can you do these by yourself and tell me what you get 4(2)-3= 4(3)-3= 4(4)-3=

OpenStudy (gamer56):

4 x 2= 8-3= 3 4 x 3=12-3= 9 4 x 4=16-3=13

OpenStudy (anonymous):

8-3=3?

OpenStudy (anonymous):

8-3=5 \[\ge\] this means a number is greater than or equal to another number 5,9,12 are greater than or equal to 5 so your solution set will be 2,3,4

OpenStudy (gamer56):

Determine which value is a solution to the open sentence. 2 + 4x = 18

OpenStudy (gamer56):

i think i got this

OpenStudy (anonymous):

subtract 2 by 18 then divide that solution by 4

OpenStudy (gamer56):

16 correct?

OpenStudy (gamer56):

wait i didnt divide

OpenStudy (anonymous):

18-2 is 16 now divide by 4

OpenStudy (gamer56):

4?

OpenStudy (anonymous):

yes

OpenStudy (gamer56):

Which is a word phrase for the variable expression? 3(n – 1)

OpenStudy (anonymous):

whats another name for multiplying

OpenStudy (gamer56):

dividing??

OpenStudy (anonymous):

no mathmatical terms for multiplying dividing would be the inverse

OpenStudy (gamer56):

oh

OpenStudy (gamer56):

so what do i do?

OpenStudy (anonymous):

well what would be a mathmatical term for multiplication?

OpenStudy (gamer56):

inverse

OpenStudy (anonymous):

no thats the opposite of a sign

OpenStudy (anonymous):

product is one mathmatical term

OpenStudy (anonymous):

and what is n and 1 doing

OpenStudy (gamer56):

subtraction

OpenStudy (gamer56):

@texaschic101 help someone

OpenStudy (anonymous):

now a mathmatical term for subtraction?

OpenStudy (gamer56):

Addition???

OpenStudy (anonymous):

no thats the inverse

OpenStudy (anonymous):

mathmatical term would be another way of saying subtraction more advance

OpenStudy (gamer56):

can you tell me?

OpenStudy (anonymous):

when you're subtracting you're finding the difference of two numbers

OpenStudy (anonymous):

soo in that equation you're finding the product of the difference of n and 1

OpenStudy (gamer56):

yes

OpenStudy (anonymous):

yes that is what the word phrase would be or similar

OpenStudy (gamer56):

So my answer would be?? @KlOwNlOvE

OpenStudy (anonymous):

well are you given options

OpenStudy (gamer56):

whoops i thought i posted them

OpenStudy (gamer56):

A. three times the quotient of a number n and 1 B. three more than the difference of a number n and 1 C. three times the difference of a number n and 1 D. one less than the product 3 and a number n

OpenStudy (anonymous):

quotient means divide and theres no dividing so A is no more means adding theres no adding so B is no C you are multiplying 3 by the difference of n and 1 D would be 3n-1

OpenStudy (anonymous):

so C is right